# Towards a Quantum Information Theory of Hadronization: Dihadron Fragmentation and Neutral Polarization in Heavy Baryons

### v1

@@ I Introduction
Every single time that quarks and gluons are produced in particle collisions, they combine back into color-neutral bound states like the proton or the pion, collectively called hadrons. Calculating this nonperturbative, strongly coupled phenomenon called hadronization from first principles using the Lagrangian of Quantum Chromodynamics (QCD) remains an open problem, restricting us to phenomenological models. Perhaps surprisingly, and despite the underlying dynamics of quark and gluons clearly being quantum mechanical (QM) in general, classical stochastic models of hadronization Field and Feynman (1978); Field (1989); Andersson et al. (1983); Webber (1984) as e.g. implemented in multi-purpose Monte-Carlo (MC) generators Bierlich et al. (2022); Bewick et al. (2024); Bothmann et al. (2024) are vastly successful in this regard, notably for open, semi-inclusive hadronization observables. A macroscopic, statistical-mechanics approach especially succeeds in the large-multiplicity environment of heavy-ion collisions Hagedorn (1965); Becattini and Heinz (1997); Andronic et al. (2003); Chojnacki et al. (2012). In this letter we aim to identify experimental measurement outcomes that indicate a departure from classical expectations and – if observed – can only be described by a full QM treatment of hadronization, see e.g. Refs. Metz (2002); Kerbizi and Lönnblad (2022, 2023); Masouminia and Richardson (2024). This, to the reader, will be a familiar task: Asking whether local reality itself is quantum in the form of Bell’s inequality Bell (1964) lies at the root of today’s rich field of quantum information (QI) theory. We will therefore adapt and extend QI concepts to hadronization physics to ask, broadly speaking: Is hadronization quantum?
@@
Our proposal is part of a larger fruitful effort to bring QI theory to bear on questions in particle physics Tornqvist (1981); Ablikim et al. (2019); Severi et al. (2022); Aoude et al. (2022); Afik and de Nova (2023); Aad et al. (2024); Barr et al. (2024); Wu et al. (2024a); Morales (2024); Chen et al. (2024); Kowalska and Sessolo (2024); Fabbrichesi and Marzola (2024); Bernal et al. (2024); Hayrapetyan et al. (2024); Afik et al. (2024); Guo et al. (2025); White and White (2024); Wu et al. (2024b); Fabbrichesi et al. (2024); Demina and Landi (2025); Ruzi et al. (2024); Morales and Szynkman (2024); Grossi et al. (2024); Cheng et al. (2024); Fabbrichesi et al. (2025a); Cheng and Yan (2025); Fabbrichesi et al. (2025b, c), but distinct in that most of the existing work, if not attempting to explain the Standard Model (SM) Mulders (2018); Thaler and Trifinopoulos (2024), aims to probe quantum foundations either at the highest energy scales or in exclusive hadronic processes, essentially asserting the quantum nature of the perturbative SM Lagrangian or hadronic form factors, respectively. Hadronization then either serves as an analyzer for the perturbative density matrix Aidala and Rogers (2021); Afik et al. (2024); Cheng and Yan (2025) or, possibly, as a source of decoherence Gong et al. (2022); Datta et al. (2024) acting on the entangled state created during the perturbative scattering. By contrast, our proposal directly concerns the nature of hadronization itself, and we appeal to principles of effective field theory and factorization theorems to isolate the nonperturbative behavior under study from the (in our view, most likely quantum, and calculable) perturbative scattering and the self-analyzing hadronic decays. Our work is also distinct from proposals to simulate hadronization on quantum circuits Florio et al. (2023); Bauer et al. (2023); Florio et al. (2024); Li et al. (2024); Lee et al. (2024); Bauer (2025).
@@ II Classical hadronization models as “hidden” variables
In the picture we are developing, Bell’s “hidden variables” $\lambda$ take on concrete meaning as a point in the phase space of a stochastic hadronization model with an intermediate classical state $p(\lambda)$, i.e., a probability density. Of course, in this case the variables are not really hidden: In an analytic model, one may calculate their statistical distribution; if the model is implicitly defined by a stochastic algorithm, e.g. a Markov process acting on an ensemble of strings or preconfined clusters Amati and Veneziano (1979); Andersson et al. (1983); Webber (1984), one can print out and bin the intermediate state to record $p(\lambda)$. Just as in Bell’s concise proof Bell and Aspect (2004) of the Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser et al. (1969), suitable correlation observables $\mathcal{C}$ will have an expectation value $\langle\mathcal{C}\rangle_{p(\lambda)}\leq\mathcal{C}_{\mathrm{max}}$ that is bounded when computed in any classical hadronization model, following basic properties of the integral measure in any number of classical phase-space dimensions, while QM expectation values can exceed the bound. Note that $p(\lambda)$ and the conditional measurement probabilities of classical hadronization models must obey the spacetime symmetries of QCD (Lorentz covariance and parity), a constraint that is typically absent in tests of quantum foundations, but which we will make productive use of in this letter.
@@
where $\hbar=1$ and $\hat{S}_{j,L}$ ($\hat{\mathbf{S}}_{j,T}$) is the longitudinal (transverse) component of the spin operator of hadron $h_{j}$ with respect to the fragmentation axis $\mathbf{z}$ pointing back to the hard collision in the hadron’s rest frame. (In the last term, a contraction between vector indices is understood.) Eq. (1) is the polarization state of the system at long distances of order the hadron lifetimes, $t_{\mathrm{decay}}\sim 1/\Gamma_{h}$, where a QM description of the system is appropriate for the exclusive hadronic decays.111 Decay matrix elements and differential distributions are collected in the supplemental material. By contrast, the nonperturbative hadronization physics are encoded in normalized dihadron fragmentation functions (DiFFs) Metz and Vossen (2016); Pitonyak et al. (2024); Rogers et al. (2025); Pitonyak et al. (2025) $\tilde{D}_{LL}$ Huang et al. (2025) and $\tilde{D}_{TT}$, which are determined from experiments or computed in a model. In the latter case, we interpret the DiFFs as matching coefficients that are fixed by computing spin expectation values $T_{k\ell}/4\equiv\langle S_{1,k}\otimes S_{2,\ell}\rangle$ in a (potentially classical) hadronization model at the scale $t_{\mathrm{confine}}\sim 1/\Lambda_{\mathrm{QCD}}\ll t_{\mathrm{decay}}$ where the hadrons form and decouple. We also have $t_{\mathrm{confine}}\gg t_{\mathrm{hard}}$, where $t_{\mathrm{hard}}$ is the very short time scale of the perturbative hard scattering. This separation of all scales allows us to specifically probe the quantum nature of hadronization and, while abstract, mirrors the workflow of MC generators.
@@
The DiFFs completely characterize the system, see Fig. 1: Positivity requires $\tilde{D}_{LL}\geq-1$ and $\tilde{D}_{LL}+2\lvert\tilde{D}_{TT}\rvert\leq 1$. By the positive-partial-transpose (PPT) criterion Horodecki et al. (1996), the hadrons are entangled if and only if $\lvert\tilde{D}_{TT}\rvert>(\tilde{D}_{LL}-1)/2$. Most interestingly, we can evaluate the optimal Horodecki et al. (1995) CHSH inequality Clauser et al. (1969) for Eq. (1) in terms of the DiFFs,
$\tilde{D}_{TT}^{2}+\max\{\tilde{D}_{LL}^{2},\tilde{D}_{TT}^{2}\}\stackrel{{\scriptstyle p(\lambda)}}{{\leq}}1\,.$ (2)
@@
This is the key result for this example process, illustrating our proposal: If Eq. (2) is found to be violated in data, the observation cannot be explained by any classical hadronization model. The proof exactly follows Ref. Bell and Aspect (2004), with one crucial difference: In our case, the “locality” assumption that measurements on $h_{1}$ are independent of the reconstruction of $h_{2}$ is guaranteed by color confinement, rather than by hypothesizing, like Bell, a local hidden-variable theory (LHVT) of Nature. The CHSH test we propose here therefore does not test quantum nonlocality in the usual way, but instead uses the identical mathematics to test whether the blackbox Hamiltonian of hadronization admits a classical description.222 Another assumption is that Gedanken measurements $\langle(\bm{e}_{\alpha}\cdot\bm{S}_{1})\otimes(\bm{S}_{2}\cdot\bm{e}_{\beta})\rangle$ at Bell angles $\alpha,\beta$ are indeed predicted to evaluate to $\bm{e}_{\alpha}\cdot(\mathsf{T}/4)\cdot\bm{e}_{\beta}$, where $\mathsf{T}=(T_{k\ell})$. In our case, this property is inherited from the interpolating partonic spin operators defined in the underlying QCD theory. Therefore, the counterexamples of Refs. Kasday et al. (1970); Abel et al. (1992) (nota bene constructed to argue against the viability of testing LHVTs of Nature at colliders) do not apply.
@@
Figure 1: QI characterization of parton fragmentation to two nearby polarized hadrons. An observation of dihadron fragmentation functions taking values in the blue region is inconsistent with classical models of hadronization. For $\tilde{D}_{LL}=-1$ and $\tilde{D}_{TT}=-1$ ($+1$), the hadrons are produced in a pure singlet (triplet) Bell state.
@@ III Quantum contextuality and local spin-1 systems
through a joint conditional measurement probability at given $\lambda$. Of particular interest are $\{\bm{e}_{i}\}$ arranged in connected Kochen-Specker (KS) graphs KOCHEN and SPECKER (1967) where $\bm{e}_{i}\perp\bm{e}_{j}$ for all edges in the graph. The first nontrivial case of a simple cyclic graph is the Klyachko-Can-Binicioglu-Shumovsky (KCBS) configuration =$\bigstar=\{\bm{e}_{i}\}$ Klyachko et al. (2008) of $N=5$ unit vectors $\bm{e}_{i}\perp\bm{e}_{i+1\!\!\!\mod 5}$ arranged around a symmetry axis $\bm{r}=\bm{z}$, i.e., they form a regular pentagram in the plane of fixed $\bm{r}\cdot\bm{e}_{i}=1/\sqrt[4]{5}$. Crucially, even though the additional measurements in Eq. (5) are counterfactual in our case, the existence of the joint distribution does leave an imprint on experimentally observable distributions. This is the statement of the KCBS inequality Klyachko et al. (2008), which in the general case constrains the sum of two-point correlations. Using $\langle\bm{J}^{2}\rangle=2$, one readily shows (also in the classical case) that Eq. (5) exists for $\bigstar$ if and only if the single-point marginal probabilities satisfy $\sum_{\bm{e}\in\bigstar}P(1|\bm{e})=\sum_{\bm{e}\in\bigstar}\langle J_{e}^{2}\rangle\geq 3$. Importantly, these latter single-point measurements are accessible experimentally and determined by $w_{1}$. We thus find a KCBS bound for the Falk-Peskin parameter, see Fig. 2,
@@
It is important to realize that classical models do not have to be noncontextual, cf. Eq. (7). Instead, the degree of contextuality is a property of the hadronization model that can be assessed by inspecting its dynamics and coupling to the final-state hadron spin. (The intuition of the KCBS-noncontextual limit being “more classical” is valuable nevertheless, cf. the case of a thermal system at high temperatures, where classical and quantum dynamics become indistinguishable and isotropic, $w_{1}\to 2/3>w_{1}^{\bigstar}$.) Conversely, the degree to which hadronization models can be coupled to hadron spin in a noncontextual way has fundamental limits. The tightest such constraint for $j=1$ arises from the Yu-Oh configuration Yu and Oh (2012), a KS graph with $N=13$ containing several cycles, whose associated state-independent noncontextuality inequality when combined with $\langle\bm{J}^{2}\rangle=2$ rules out the existence of Eq. (5) for this case altogether.
@@ V Bipartite QM models and factorization
We conclude our analysis of Eq. (3) by interpreting it in a simple QM model of hadronization where the hadron is formed from the heavy quark and two light spin-$1/2$ degrees of freedom, which we can e.g. think of as constituent quarks. In this case it is interesting to ask whether the state of the light system is entangled or separable when the hadron is formed. To assess this, we convert Eq. (3) to a two-qubit density matrix (noting that the singlet state, which would correspond to the $\Lambda_{Q}$ baryon, is not populated at this binding energy), and again evaluate the PPT criterion Horodecki et al. (1996). The result for the lowest eigenvalue of the partial transpose is shown in red in Fig. 2: As expected, an overlap of $\operatorname{Tr}[\rho\,\lvert 0_{z}\rangle\langle 0_{z}\rvert]=1-w_{1}>1/2$ with $\lvert 0_{z}\rangle$, which acquires an interpretation as a triplet Bell state in this model, implies entanglement (and eventually leads to KCBS violation Klyachko et al. (2008)).
@@
where $\Phi_{q,\bar{q}}$ are (anti)quark correlation functions of the (possibly polarized) incoming protons, $\operatorname{tr}$ denotes a trace over Dirac indices, and $\mu,\nu$ encode the polarization of the virtual photon. Eq. (10) is manifestly separable with respect to the product $\mathcal{H}_{q}\otimes\mathcal{H}_{\bar{q}}$ of the (anti)quark Hilbert spaces. Conversely, the neutral polarization state $\tfrac{1}{\sqrt{2}}\bigl{(}\lvert\uparrow_{q}\downarrow_{\bar{q}}\rangle+\lvert\downarrow_{q}\uparrow_{\bar{q}}\rangle\bigr{)}$, which is maximally entangled, cannot have large overlap with Eq. (10), and indeed it is well known Levelt and Mulders (1994); Boer (2009); Bacchetta et al. (2019); Vladimirov et al. (2022); Ebert et al. (2022); Gamberg et al. (2022); del Castillo et al. (2024); Gao et al. (2025) that neutral photon polarization requires an additional spin-$1$ degree of freedom (and thus subleading-power factorization). The above argument provides a novel perspective on this fact from QI theory.

### v2

@@ I Introduction
Every single time that quarks and gluons are produced in particle collisions, they combine back into color-neutral bound states like the proton or the pion, collectively called hadrons. Calculating this nonperturbative, strongly coupled phenomenon called hadronization from first principles using the Lagrangian of Quantum Chromodynamics (QCD) remains an open problem, restricting us to phenomenological models. Perhaps surprisingly, and despite the underlying dynamics of quarks and gluons clearly being quantum mechanical (QM) in general, classical stochastic models of hadronization Field and Feynman (1978); Field (1989); Andersson et al. (1983); Webber (1984) as e.g. implemented in multi-purpose Monte-Carlo (MC) generators Bierlich et al. (2022); Bewick et al. (2024); Bothmann et al. (2024) are vastly successful in this regard, notably for open, semi-inclusive hadronization observables. A macroscopic, statistical-mechanics approach especially succeeds in the large-multiplicity environment of heavy-ion collisions Hagedorn (1965); Becattini and Heinz (1997); Andronic et al. (2003); Chojnacki et al. (2012). In this letter we aim to identify experimental measurement outcomes that indicate a departure from classical expectations and – if observed – can only be described by a full QM treatment of hadronization, see e.g. Refs. Metz (2002); Kerbizi and Lönnblad (2022, 2023); Masouminia and Richardson (2024). This, to the reader, will be a familiar task: Asking whether local reality itself is quantum in the form of Bell’s inequality Bell (1964) lies at the root of today’s rich field of quantum information (QI) theory. We will therefore adapt and extend QI concepts to hadronization physics to ask, broadly speaking: Is hadronization quantum?
@@
Our proposal is part of a larger fruitful effort to bring QI theory to bear on questions in particle physics Tornqvist (1981); Ablikim et al. (2019); Neill and Waalewijn (2019); Afik and de Nova (2021); Takubo et al. (2021); Severi et al. (2022); Afik and de Nova (2022); Aoude et al. (2022); Afik and de Nova (2023); Dong et al. (2024); Morales (2023); Aad et al. (2024); Barr et al. (2024); Wu et al. (2024a); Morales (2024); Chen et al. (2024); Kowalska and Sessolo (2024); Fabbrichesi and Marzola (2024); Bernal et al. (2024); Hayrapetyan et al. (2024); Afik et al. (2025); Guo et al. (2025); White and White (2024); Wu et al. (2024b); Fabbrichesi et al. (2024); Demina and Landi (2025); Lo Chiatto (2024); Ruzi et al. (2024); Morales and Szynkman (2024); Grossi et al. (2024); Cheng et al. (2024); Fabbrichesi et al. (2025a); Cheng and Yan (2025); Fabbrichesi et al. (2025b, c), but distinct in that most of the existing work, if not attempting to explain the Standard Model (SM) Cervera-Lierta et al. (2017); Mulders (2018); Thaler and Trifinopoulos (2025), aims to probe quantum foundations either at the highest energy scales or in exclusive hadronic processes, essentially asserting the quantum nature of the perturbative SM Lagrangian or hadronic form factors, respectively. Hadronization then either serves as an analyzer for the perturbative density matrix Aidala and Rogers (2021); Afik et al. (2025); Cheng and Yan (2025) or, possibly, as a source of decoherence Gong et al. (2022); Datta et al. (2025) acting on the entangled state created during the perturbative scattering. By contrast, our proposal directly concerns the nature of hadronization itself, and we appeal to principles of effective field theory and factorization theorems to isolate the nonperturbative behavior under study from the (in our view, most likely quantum, and calculable) perturbative scattering and the self-analyzing hadronic decays. Our work is also distinct from proposals to simulate hadronization on quantum circuits Florio et al. (2023); Bauer et al. (2023); Florio et al. (2024); Li et al. (2024a); Lee et al. (2025a); Bauer (2025).
@@ II Classical hadronization models as “hidden” variables
In the picture we are developing, Bell’s “hidden variables” $\lambda$ take on concrete meaning as a point in the phase space of a stochastic hadronization model with an intermediate classical state $p(\lambda)$, i.e., a probability density. Of course, in this case the variables are not really hidden: In an analytic model, one may calculate their statistical distribution; if the model is implicitly defined by a stochastic algorithm, e.g. a Markov process acting on an ensemble of strings or preconfined clusters Amati and Veneziano (1979); Andersson et al. (1983); Webber (1984), one can print out and bin the intermediate state to record $p(\lambda)$. One then shows that for large classes of such models, suitable correlation observables $\mathcal{C}$ have a bounded expectation value $\langle\mathcal{C}\rangle_{p(\lambda)}\leq\mathcal{C}_{\mathrm{max}}$, using basic properties of the integral measure in any number of classical phase-space dimensions, while QM expectation values can exceed the bound. Note that $p(\lambda)$ and the conditional measurement probabilities of classical hadronization models must obey the spacetime symmetries of QCD (Lorentz covariance and parity), a constraint that is typically absent in tests of quantum foundations, but which we will make productive use of in this letter.
@@
where $\hbar=1$ and $\hat{S}_{j,L}$ ($\hat{\mathbf{S}}_{j,T}$) is the longitudinal (transverse) component of the spin operator of hadron $h_{j}$ with respect to the fragmentation axis $\mathbf{z}$ pointing back to the hard collision in the hadron’s rest frame. (In the last term, a contraction between vector indices is understood.) Eq. (1) is the polarization state of the system at long distances of order the hadron lifetimes, $t_{\mathrm{decay}}\sim 1/\Gamma_{h}$, where a QM description of the system is appropriate for the exclusive hadronic decays sup . By contrast, the nonperturbative hadronization physics are encoded in normalized dihadron fragmentation functions (DiFFs) Metz and Vossen (2016); Pitonyak et al. (2024); Rogers et al. (2025); Pitonyak et al. (2025) $\tilde{D}_{LL}$ Huang et al. (2025) and $\tilde{D}_{TT}$, which are determined from experiments or computed in a model. In the latter case, we interpret the DiFFs as matching coefficients that are fixed by computing spin expectation values $T_{k\ell}/4\equiv\langle S_{1,k}\otimes S_{2,\ell}\rangle$ in a (potentially classical) hadronization model at the scale $t_{\mathrm{confine}}\sim 1/\Lambda_{\mathrm{QCD}}\ll t_{\mathrm{decay}}$ where the hadrons form and decouple. We also have $t_{\mathrm{confine}}\gg t_{\mathrm{hard}}$, where $t_{\mathrm{hard}}$ is the very short time scale of the perturbative hard scattering. This separation of all scales allows us to specifically probe the quantum nature of hadronization and, while abstract, mirrors the workflow of MC generators.
@@
The DiFFs completely characterize the system, see Fig. 1: Positivity requires $\tilde{D}_{LL}\geq-1$ and $\tilde{D}_{LL}+2\lvert\tilde{D}_{TT}\rvert\leq 1$. By the positive-partial-transpose (PPT) criterion Horodecki et al. (1996), the hadrons are entangled if and only if $\lvert\tilde{D}_{TT}\rvert>(\tilde{D}_{LL}-1)/2$. Most interestingly, we can consider the class of fully classical hadronization models where the (spin) properties of individual final-state hadrons are separately determined by a common underlying classical state $\lambda$,
$\langle S_{1,k}\otimes S_{2,\ell}\rangle=\int\!\mathrm{d}\lambda\,p(\lambda)\,S_{1,k}(\lambda)\,S_{2,\ell}(\lambda)\,.$ (2)
For these models, the following optimal Horodecki et al. (1995) Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser et al. (1969) holds,
[…]
@@
This is the key result for this example process, illustrating our proposal: If Eq. (3) is found to be violated in data, the observation cannot be explained by any fully classical hadronization model of the form Eq. (2).111 By comparison, the observed Bose-Einstein enhancement of nearby identical hadrons Andersson and Hofmann (1986); Artru and Bowler (1988); Acton et al. (1993) simply follows from (anti)symmetrizing the wave function in momentum space. Doing so on top of a classical model leaves the normalized Eq. (1) unaffected, making Eq. (3) a deeper test of genuine QM dynamics. The derivation exactly follows Bell’s concise proof of the CHSH inequality in Ref. Bell and Aspect (2004), with one crucial difference: In our case, the “locality” assumption in Eq. (2) is a property that can be checked for any given hadronization model (without computation) by inspecting its dynamical rules and coupling to hadron spins, and not a property of a Bell-type local hidden-variable theory (LHVT) of Nature. The CHSH test we propose here therefore does not test quantum nonlocality in the usual way, but instead uses the identical mathematics to test whether the blackbox Hamiltonian of hadronization admits a classical description.222 Another assumption is that Gedanken measurements $\langle(\bm{e}_{\alpha}\cdot\bm{S}_{1})\otimes(\bm{S}_{2}\cdot\bm{e}_{\beta})\rangle$ at Bell angles $\alpha,\beta$ are indeed predicted to evaluate to $\bm{e}_{\alpha}\cdot(\mathsf{T}/4)\cdot\bm{e}_{\beta}$, where $\mathsf{T}=(T_{k\ell})$. In our case, this property is inherited from the interpolating partonic spin operators defined in the underlying QCD theory. Therefore, the counterexamples of Refs. Kasday et al. (1970); Abel et al. (1992); Li et al. (2024b) (nota bene constructed to argue against the viability of testing LHVTs of Nature at colliders) do not apply.
@@
Figure 1: QI characterization of parton fragmentation to two nearby polarized hadrons. An observation of dihadron fragmentation functions taking values in the blue region is inconsistent with classical models of hadronization. For $\tilde{D}_{LL}=-1$ and $\tilde{D}_{TT}=-1$ ($+1$), the hadrons are produced in a pure singlet (triplet) Bell state. For the experimental data, see Ref. STAR Collaboration (2025); sup .
@@ III Quantum contextuality and local spin-1 systems
through a joint conditional measurement probability at given $\lambda$. Of particular interest are $\{\bm{e}_{i}\}$ arranged in connected Kochen-Specker (KS) graphs Kochen and Specker (1967) where $\bm{e}_{i}\perp\bm{e}_{j}$ for all edges in the graph. The first nontrivial case of a simple cyclic graph is the Klyachko-Can-Binicioglu-Shumovsky (KCBS) configuration =$\bigstar=\{\bm{e}_{i}\}$ Klyachko et al. (2008) of $N=5$ unit vectors $\bm{e}_{i}\perp\bm{e}_{i+1\!\!\!\mod 5}$ arranged around a symmetry axis $\bm{r}=\bm{z}$, i.e., they form a regular pentagram in the plane of fixed $\bm{r}\cdot\bm{e}_{i}=1/\sqrt[4]{5}$. Crucially, even though the additional measurements in Eq. (6) are counterfactual in our case, the existence of the joint distribution does leave an imprint on experimentally observable distributions. This is the statement of the KCBS inequality Klyachko et al. (2008), which involves two-point correlations. Using $\langle\bm{J}^{2}\rangle=2$, one readily shows (also in the classical case) that Eq. (6) exists for $\bigstar$ if and only if the single-point marginal probabilities satisfy $\sum_{\bm{e}\in\bigstar}\langle J_{e}^{2}\rangle\geq 3$. Importantly, these latter single-point measurements are accessible experimentally and determined by $w_{1}$. We thus find a KCBS bound for the Falk-Peskin parameter, see Fig. 2,
@@
It is important to realize that classical models do not have to be noncontextual, cf. Eq. (8). Instead, the degree of contextuality, like Eq. (2), is a property of the hadronization model that can be assessed by inspecting its dynamics and coupling to the final-state hadron spin. (The intuition of the KCBS-noncontextual limit being “more classical” is valuable nevertheless, cf. the case of a thermal system at high temperatures, where classical and quantum dynamics become indistinguishable and isotropic, $w_{1}\to 2/3>w_{1}^{\bigstar}$.) Conversely, the degree to which hadronization models can be coupled to hadron spin in a noncontextual way has fundamental limits. The tightest such constraint for $j=1$ arises from the Yu-Oh configuration Yu and Oh (2012), a KS graph with $N=13$ containing several cycles, whose associated state-independent noncontextuality inequality when combined with $\langle\bm{J}^{2}\rangle=2$ rules out the existence of Eq. (6) for this case altogether.
@@ V Bipartite QM models and factorization
Lastly, we interpret Eq. (4) in a simple QM model of hadronization where the hadron is formed from the heavy quark and two light spin-$1/2$ degrees of freedom, e.g. constituent quarks. Here it is interesting to ask whether the light system is entangled or separable when the hadron is formed. To assess this, we convert Eq. (4) to a two-qubit density matrix (noting that the singlet state, which would correspond to the $\Lambda_{Q}$ baryon, is not populated at this binding energy), and again evaluate the PPT criterion Horodecki et al. (1996). The result for the lowest eigenvalue of the partial transpose is shown in red in Fig. 2: As expected, an overlap of $\operatorname{Tr}[\rho\,\lvert 0_{z}\rangle\langle 0_{z}\rvert]=1-w_{1}>1/2$ with $\lvert 0_{z}\rangle$, which acquires an interpretation as a triplet Bell state in this model, implies entanglement (and eventually leads to KCBS violation Klyachko et al. (2008)).
@@
where $\Phi_{q,\bar{q}}$ are (anti)quark correlation functions of the (possibly polarized) incoming protons, $\operatorname{tr}$ denotes a trace over Dirac indices, and $\mu,\nu$ encode the polarization of the virtual photon. Eq. (11) is manifestly separable with respect to the product $\mathcal{H}_{q}\otimes\mathcal{H}_{\bar{q}}$ of the (anti)quark Hilbert spaces Boer et al. (2005). Conversely, the neutral polarization state $\tfrac{1}{\sqrt{2}}\bigl{(}\lvert\uparrow_{q}\downarrow_{\bar{q}}\rangle+\lvert\downarrow_{q}\uparrow_{\bar{q}}\rangle\bigr{)}$, which is maximally entangled, cannot have large overlap with Eq. (11), and indeed it is well known Levelt and Mulders (1994); Boer (2009); Bacchetta et al. (2019); Vladimirov et al. (2022); Ebert et al. (2022); Gamberg et al. (2022); del Castillo et al. (2024); Gao et al. (2025) that neutral photon polarization requires an additional spin-$1$ degree of freedom (and thus subleading-power factorization). The above argument provides a novel perspective on this fact from QI theory.